OpenStudy (anonymous):

A box contains 15 balls numbered 1 through 15. Two balls are drawn in succession without replacement. If the second ball has the number 4 on it, What is the probability that the first ball had a smaller number on it? An even number on it?

OpenStudy (anonymous):

Of all the balls there are only 3 with smaller number than 4. That gives you a probability of picking one of those 3 out of the 15 as: \[P(ball\leq 3)=\frac{3}{15}\] By the same principle there are 7 even numbers \[2, 4, 6, 8, 10, 12, 14\] Probability of picking an even numbered ball is therefor \[P(even)=\frac{7}{15}\]

OpenStudy (anonymous):

The chance of it being a smaller number is 1/5. The chance of it being even is 7/15.

OpenStudy (anonymous):

Reason: 3 of the balls have smaller numbers than 4, and there are 15 balls, so the chance is 3/15, and if you divide the top and bottom of this fraction by 3 you get 1/5. There are seven even numbered balls and fifteen balls in all, so the chance of the second part is 7/15.

OpenStudy (anonymous):

Please click that medal next to my answer.

OpenStudy (anonymous):

the answer is not correct.. the answer was 3/14 and 3/7